can someone explain on how to do this problem?

can someone explain on how to do this problem?Not

Not really sure where to start on how to solve
Answer #1

If there is enough force to cause sliding on all interfaces, (we will check that there is) the tension in the rope will equal the friction force attempting to move the smaller block away from the wall

the normal force of the lower block on the upper block is the weight of the upper block
The friction force between the two blocks is the coefficient of friction times the normal force.

T = Ff = μmg = 0.2*5*9.8= 9.8 N

to check that there is enough force to cause motion between all touching surfaces, we add the tension force to the friction force with the ground and expect it to be less than the applied force

F =0.8 + 0.2(5 + 10)(9.8) = 39.2 N

as this is less than the applied force of 50 N, the Tension value is valid.

therefore answer is B)